mathslover (mathslover):

Question - Coordinate Geometry (Circles) I have been studying circles and came with this in my book : The equation of the chord of the circle \(x^2 + y^2 = a^2\) joining \((a\cos \alpha , a\sin \alpha) \) and \((a\cos \beta, a\sin \beta)\) is : \(x \cos (\dfrac{\alpha + \beta}{2}) + y\sin (\dfrac{\alpha + \beta}{2}) = a \cos (\dfrac{\alpha - \beta}{2}) \)

mathslover (mathslover):

How can I prove the above equation?

OpenStudy (accessdenied):

We might start by finding the slope of the line joining those two points and setting up an equation using point-slope form: y - y0 = m (x - x0) It probably won't look very pretty yet so we would need to do some hairy simplification work to get to that equation.

mathslover (mathslover):

\(a\sin \beta - a\sin \alpha = m ( a\cos \beta - a \cos \alpha) \\ m = \dfrac{a\sin \beta - a\sin \alpha }{a\cos \beta - a \cos \alpha} \)

OpenStudy (accessdenied):

Oh, we only use one of the points in point-slope. We want to leave the plain x and y so that in the end we have an equation with x and y. \( y - \color{blue}{y_0} = m \left( x - \color{blue}{x_0} \right) \)

mathslover (mathslover):

Oh okay, fine! So, shall I take (\(x_0 , y_0\) ) \(\equiv\) (\(a \cos \alpha , a\sin \alpha\)) or (\(a \cos \beta , a \sin \beta\) )

OpenStudy (accessdenied):

I believe both will work out in the end, so either choice should be fine. I'd go with the first one.

mathslover (mathslover):

Okay.. Great! I will work with your choice.